OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-2,-1) and (3,5).

OpenStudy (anonymous):

Y=6/5X- that was my answer and it says i got it wrong @Nnesha

Nnesha (nnesha):

you have to write whole equation y=mx + b where is your b

OpenStudy (anonymous):

use y-y1=m(x-x1) where y1 is the y value of one of our points and x1 is the x value of the same point. m will be the slope. to find the slope use y-y1/x-x1

OpenStudy (anonymous):

y+1/x+2

OpenStudy (anonymous):

you use both points to find the slope

Nnesha (nnesha):

slope formula is